One number is 4 times a first number. A third number is 100 more than the first number. If the sum of the three numbers is 286, find the number.

Answer by checkley79(3341)

LET X BE THE FIRST NUMBER.
4X IS THE NEXT NUMBER.
X=100 IS THE NEXT NUMBER.
X+4X+X+100=286
6X=286-100
6X=186
X=186/6
X=31 ANS. FOR THE FIRST NUMBER.
4*31=124 ANS. FOR THE NEXT NUMBER.
31+100=131 ANS. FOR THE NEXT NUMBER.
PROOF:
31+124+131=286
286=286